OpenStudy (anonymous):

Which of the following is the conjugate of a complex number with 6 as the real part and -5 as the imaginary part? 6 - 5i -6 - 5i 6 + 5i -6 + 5i

OpenStudy (kmeis002):

give a complex number of the form \[z = a+bi \] We label the complex conjugate by: \[\bar{z}=a-bi \] . Example. If z = 2 - 4i, find its complex conjugate. Ans:\[\bar{z} = 2+4i \]